OpenStudy (anonymous):

A long air-cored solenoid has two windings wound on top of each other. Each has N turns per metre and resistance R. Deduce expressions for the flux density at the centre of the solenoid when the windings are connected A)In series B)In parallel To the battery of emf E and of negligible internal resistance (in each case the magnetic field produced by the currents in the two windings reinforce.) YOU CAN REFER ROGER MUNCASTER page 728 F99

OpenStudy (irishboy123):

all you're doin' here is using \(\large B = \mu_0 \frac{N I}{l}\) to establish the magnetic flux density , but first you need I, the current, so Ohm's law for resistors, R and R, in series and then in parallel

OpenStudy (anonymous):

I=E/(R+R)=E/(2R) series connection I=E/(1/R+1/R)=E/(2/R) for parallel connection

OpenStudy (irishboy123):

just check your second one. \(R_T = \dfrac{R}{2}\) so I = E/(R/2)

OpenStudy (anonymous):

Correct it is I=E/(R/2) I didn't reciprocate... Thanks much

OpenStudy (radar):

Was just wondering if the "sense" of the magnetic field would need to be considered. It is possible to connect in series with the turns direction in the same or in the opposite direction